OpenStudy (anonymous):

Describe the nature of the roots for this equation. x^2+3x-7=0 A.Two complex roots B.One real, double root C.Two real, rational roots D.Two real, irrational roots

OpenStudy (anonymous):

The trick to this is to use the descriminant of the quadratic.

OpenStudy (anonymous):

Are you familiar with the term?

OpenStudy (anonymous):

determinant*

OpenStudy (anonymous):

oops discriminant*

OpenStudy (anonymous):

delta = b^2 - 4(a*c)

OpenStudy (anonymous):

ya

OpenStudy (anonymous):

Ok: \[b^2 - 4ac > 0 \implies \text{ 2 real roots}\] \[b^2 - 4ac < 0 \implies \text{ 2 imaginary(complex) roots}\] \[b^2 - 4ac = 0 \implies \text{ 1 real root}\] And if \(\sqrt{b^2 - 4ac}\) is irrational , the roots will be irrational.

OpenStudy (anonymous):

so d?